OpenStudy (tabithax3):

Find the foci of the ellipse: 2x^2+y^2=3

OpenStudy (jdoe0001):

do you know how to put the equation in ellipse form?

OpenStudy (tabithax3):

@jdoe0001 i got \[\frac{ 2x ^{2} }{ 3 }+\frac{ y^2 }{ 3 }=1\] not sure if i did that right...

OpenStudy (jdoe0001):

I got the same :D

OpenStudy (jdoe0001):

or \[\frac{x^2}{\frac{3}{2}}+\frac{y^2}{3}\]

OpenStudy (tabithax3):

@jdoe0001 how did you change \[x^{2}\]?

OpenStudy (jdoe0001):

\[\frac{2x^2}{3}=\frac{\alpha}{\beta}*X^2 = \frac{x^2}{\frac{\beta}{\alpha}}\]

OpenStudy (jdoe0001):

if you had a fraction as denominator, you simply flip it and multiply against the numerator to get a 'whole' fraction

OpenStudy (jdoe0001):

\[so\ your\ a=\sqrt{\frac{3}{2}}\ and\ your\ b= \sqrt{3}\] becuause both the 'a' and 'b' on the ellipse have to be squared, so \[so\ your\ a^2=({\sqrt{\frac{3}{2}}})^2\ and\ your\ b^2= ({\sqrt{3}})^2\]

OpenStudy (tabithax3):

@jdoe0001 thank you so much!

OpenStudy (jdoe0001):

yw